Solve for the indicated variable. Include all of your work in your answer. Submit your solution.
C=5(F-32)/9;for F

Answer by ewatrrr(24785)

C=5(F-32)/9
9C=5(F-32)
(9/5)C=(F-32)
(9/5)C + 32 = F